OpenStudy (anonymous):

simplify 5 over 4 + 3i and express in a + bi form. I already took the conjugate and got 20-5i/25. how do I express in a+bi form? am I finished?

OpenStudy (misty1212):

\[\frac{20-5i}{25}=\frac{20}{25}-\frac{5}{25}i\] but you probably have to reduce the fractions

OpenStudy (anonymous):

ohhh I always forget about that step! thank you so much

OpenStudy (misty1212):

is it \[\frac{20-5i}{25}=\frac{20}{25}-\frac{5}{25}i\] or \[\frac{20-15i}{25}=\frac{20}{25}-\frac{15}{25}i\]

OpenStudy (misty1212):

maybe the second one because you have to distribute the 5 when you multiply

OpenStudy (anonymous):

15 whoops

OpenStudy (anonymous):

so 4/5 - 3i/5 thanks so much for the help

OpenStudy (misty1212):

you're welcome